OpenStudy (meilendurcer):

Match each expression to its value when a=3 and b=5.

OpenStudy (meilendurcer):

\[\sqrt{3^{2}+3\times5^{2}-3}\] \[-\sqrt{20\times5}+3\times3^{2}\] \[5^{2}-\sqrt{3^{2}\times4^{2}}\] \[3^{2}-\sqrt{11+5^{2}}\]